OpenStudy (anonymous):

Evaluate each function at the given value of x. 1. g(x) = -4x^2+ 6 when x = 3 2. f(x) = 6x^4 - x^3 + 3x^2 - 5x + 9 when x = -1 3. g(x) = x^3 - 8x + 6 when x = -3.

OpenStudy (anonymous):

Simply substitute 1) x=3 into y= -4x^2 + 6 2) x=-1 into y = 6x^4 -x^3 + 3x^2 - 5x + 9 3) x=-3 into y=x^3 -8x + 6

OpenStudy (anonymous):

ok say im doing number 1, I would take the 3 and put it where the (x) is?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

ok, thanks.